### v1

@@ 2.2.4 Environmental sustainability
In AI regulation, the concept of environmental sustainability emphasizes the creation, implementation, and utilization of AI systems in ways that mitigate adverse environmental effects. This approach recognizes the ecological impact of AI technology, focusing on issues such as energy use, efficient resource management, and promoting ecological sustainability [15]. This includes safeguarding the environment and enhancing its quality, which also pertains to human health and safety [21].
Crucial aspects of this environmental focus in AI regulation include promoting energy efficiency, particularly in AI systems with high computational demands like large data centers, which are significant electricity consumers [28]. Additionally, it encompasses the pursuit of sustainable development, efforts to decrease the carbon footprint of AI technologies, and the advancement of ‘Green AI’ [14]. This latter term refers to designing AI models that are not just high-performing but also optimized for minimal consumption of computational resources and energy. Moreover, it involves adopting eco-friendly practices, such as minimizing electronic waste and encouraging the recycling of AI hardware components.
@@ 2.2.5 Traceability
The principle of traceability within the framework of AI regulation encompasses the systematic capability to monitor, record, and chronicle the decision-making processes employed by AI systems during their entire operational lifespan [21]. This principle is crucial for ensuring that AI operations are transparent and accountable. For instance, an AI system should possess robust logging features that provide a traceability level suitable for its intended use, as per the guidelines outlined in the AI Act. Such traceability measures are vital for verifying the AI system’s performance and compliance throughout its lifecycle [42].
Additionally, traceability involves closely tracking the various iterations of AI models, encompassing all modifications and updates that occur over time. This tracking is not just a record-keeping exercise but a critical tool for understanding how an AI model evolves [43]. It enables stakeholders to comprehend the changes made to the model, assess their impacts, and, if necessary, revert to prior versions when newer updates do not perform as expected or introduce unforeseen issues.
@@ 2.3 Human-centered AI Act
We categorize stakeholders into four main groups: enterprises, researchers and developers, regulators and policymakers, and users and customers, as shown in Figure 3. According to recent research [44, 45, 46, 47], these stakeholders have distinct interests and responsibilities regarding ethical AI development and deployment. Enterprises are responding to both internal and external pressures to ensure their AI technologies are ethically deployed and do not cause harm. For instance, Microsoft conducted a Human Rights Impact Assessment (HRIA) for AI, and Google has established AI principles that reference human rights [44, 45]. These efforts aim to mitigate adverse impacts, ensure corporate social responsibility, and maintain public trust. Regulators and policymakers are responsible for aligning AI with human rights principles throughout each stage of the AI lifecycle. Researchers and developers, who are deeply involved in each lifecycle stage, ensure technical performance and efficiency while promoting responsible AI behavior. They also contribute to the discourse on AI and human rights by examining the social impacts of AI, developing ethical frameworks, and advocating for policies that prioritize human dignity [45]. Lastly, end-users and customers focus on the risks and harms associated with AI. They advocate for AI systems that are free from biases, transparent in their operations and decision-making processes, particularly concerning conflicts of interest, and safe and traceable, including maintaining data integrity and security [46].
@@
Ethical concerns in the data preparation stage include ensuring data safety, obtaining informed consent, and addressing biases in data collection [48]. It is crucial to collect representative data that accurately reflects the diversity of the population to avoid discrimination and ensure fairness. This process must also ensure the data owners’ safety and make the data collection process transparent and traceable for data providers. During model training, it is essential to build a robust AI against various threats to ensure its safety, while the ethical use of computational resources should be considered to minimize environmental impacts. Transparency and traceability in the deployment process of AI models, including integrating other components within the AI system, are crucial to ensuring user safety. During the inference stage, it is important to evaluate the model’s fairness, environmental sustainability, and safety in real-world scenarios [48], and accordingly can opt to implement bias mitigation, cost-effective, and threats mitigation techniques.
@@ 3.1.1 Overview
There’s a concerted effort within the AI security community to compile and comprehend the TTPs that adversaries might deploy against AI systems. A notable contribution in this domain comes from MITRE [49], renowned for their MITRE ATT&CK TTP framework. They have extended their expertise to the AI field by publishing a comprehensive set of TTPs tailored for AI systems [50]. This work by MITRE is instrumental in aiding organizations to better understand and prepare for potential threats against their AI infrastructures, reflecting the growing importance of specialized TTPs in the evolving landscape of AI security.
@@
Table 1: Potential Attacks Across Different Stages of the AI Pipeline with a Focus on Ensuring Safety
@@
Data Poisoning Deliberate manipulation of training data to corrupt model learning, causing incorrect predictions. Common methods include injecting malicious data or altering existing data points. Data Acquisition and Preparation [51][52]
Training Data Extraction Unauthorized extraction or reconstruction of sensitive information from the training dataset used in model development. Model Deployment and Inference [53]
Training-time Attacks in GenAI Attacks targeting pre-training and fine-tuning stages of GenAI models by manipulating portions of training data to cause specific model failures. All stages [54]
Backdoor Attack Insertion of hidden triggers into AI models that cause incorrect behavior only when specific inputs are present, while maintaining normal performance otherwise. All stages [55][56]
Adversarial Examples Attack Specially crafted inputs that cause AI models to produce incorrect outputs despite appearing normal to humans. Includes white-box, black-box, and transferability attacks. Model Deployment and Inference [57][58][59]
Exfiltration Theft of sensitive data, proprietary algorithms, or intellectual property from AI models through unauthorized access or exploitation. Model Deployment and Inference [60][61]
Inference-time attacks in GenAI Exploitation of vulnerabilities in deployed LLMs and RAG applications during the model’s operational phase. Model Deployment and Inference [62]
Prompt Attacks Manipulation of LLM inputs through carefully crafted prompts to exploit model vulnerabilities and induce unintended behaviors. All stages [63]
Privacy Attacks Unauthorized extraction of sensitive information through techniques like membership inference, model extraction, and property inference attacks. All stages [64]
Ethical Hacking Authorized security testing of AI systems to identify and address vulnerabilities before malicious exploitation. All stages [65]
@@ 3.1.2 Techniques related to safety and their relations to other aspects of the AI Act
Adversarial training: Adversarial training, as introduced by Goodfellow et al. [37] and further developed by Madry et al. [66], serves as a compelling example of this principle in action. This method involves augmenting training data with adversarial examples, which are crafted iteratively during the training process using their correct labels. As the adversarial attacks used to generate these examples grow stronger, the resulting model becomes more resilient to similar types of attacks, effectively boosting the system’s safety.
One of the key benefits of adversarial training is that it leads to models that encapsulate more semantic meaning compared to standard models [67]. This aspect is particularly relevant to the concept of explainability. Models with higher semantic content are generally more interpretable, meaning their behavior can be understood and anticipated more effectively by humans. This understanding is crucial for safety, as it allows developers and users to predict and mitigate potential failures or adversarial exploits. However, integrating adversarial training into AI systems does come with trade-offs. Notably, there’s typically a decrease in model accuracy on clean, non-adversarial data. This reduction in performance highlights the delicate balance between robustness (i.e., resilience to adversarial attacks) and accuracy in typical machine learning applications. Furthermore, adversarial training is computationally expensive, primarily due to the iterative generation of adversarial examples during the training phase. This complexity underscores the resource-intensive nature of developing explainable and robust AI systems.
@@
Formal verification employs methods from formal methods to certify the robustness of neural networks. Techniques like Satisfiability Modulo Theories (SMT) solvers and abstract interpretation are used to verify the robustness of neural networks against adversarial inputs [68] [69]. Formal verification provides a rigorous, mathematically grounded process for verifying the robustness of AI systems. This rigorousness is invaluable for transparency, as it provides a solid foundation for understanding and documenting the model’s behavior under adversarial conditions. Despite current limitations (e.g., lack of scalability, computational cost, etc.), the potential of formal verification techniques is significant. They offer the possibility of comprehensive, mathematically rigorous certification of neural network robustness, which is a cornerstone for transparency. Understanding precisely how and why a model is robust against certain types of adversarial inputs allows for a more transparent AI system.
Certified adversarial robustness: Certified adversarial robustness is crucial for enhancing AI system safety, ensuring that AI models behave predictably and reliably even when faced with adversarial inputs. This robustness is especially significant from the perspective of transparency, which involves understanding and tracking the behavior and performance of AI systems under various conditions.
@@
Randomized smoothing is a technique that enhances the robustness of classifiers by integrating Gaussian noise into the input data [70][71]. The core idea is to generate predictions that are most likely correct under these noise perturbations. This approach yields provably robust classifiers, particularly against $\ell_{2}$-norm [53]. By ensuring that the classifier provides the most likely prediction under noise, randomized smoothing enhances the predictability of the AI system. This predictability is crucial for transparency, as it ensures that the system’s behavior remains consistent and understandable, even in the presence of slight input variations. The method offers certified robustness for a subset of testing samples, depending on specific factors like the radius of the $\ell_{2}$ ball and the dataset’s characteristics. This certification provides tangible evidence of the model’s robustness, aiding in transparency by offering clear, quantifiable metrics to track and understand the model’s performance under adversarial conditions. Recent advancements have extended this concept to more complex perturbations, integrating denoising diffusion probabilistic models with high-accuracy classifiers. This extension broadens the scope of transparency, allowing stakeholders to understand and certify the model’s behavior under a wider array of adversarial scenarios[53].
Training data sanitization: Training data sanitization plays a crucial role in enhancing the safety of AI systems by mitigating the impact of poisoned samples. These methods are rooted in the understanding that poisoned samples often exhibit distinct characteristics compared to regular training samples. Data sanitization techniques are purposefully designed to cleanse the training dataset of these malicious samples before machine learning training takes place. By actively excluding poisoned samples from the training sample set, techniques like the Region of Non-Interest (RONI) method enhance the model’s reliability [72]. Improved sanitization methods focus on reducing computational complexity while maintaining the robustness of the cleaning process. For instance, label cleaning methods target label flipping attacks specifically, ensuring that the model’s learning process isn’t misled by incorrectly labeled data [73]. Outlier detection [74] and clustering methods [75][76] further refine the dataset by identifying and removing anomalies, making the training data more representative of genuine patterns. Additionally, the practice of computing variance in predictions across an ensemble of models aids in recognizing inconsistencies, potentially flagged as data poisoning attempts[77]. Ensuring dataset protection through cybersecurity mechanisms for provenance and integrity attestation post-sanitization adds another layer of security. Altogether, training data sanitization serves as a critical safeguard, making AI systems more transparent by removing malicious samples and preserving the authenticity of the training data, ultimately enhancing the system’s safety and decision-making reliability.
@@
Robust training: Robust training enhances the transparency and safety of AI systems by integrating techniques like ensemble model voting, robust optimization, and randomized smoothing [78][79]. Ensemble models allow for comparative analysis of individual model decisions, improving insight into the decision-making process. Robust optimization, through methods like trimmed loss functions, minimizes the impact of extreme data points, making the model’s reasoning more transparency and less prone to data poisoning [80][81]. Randomized smoothing adds stability to model predictions against minor input perturbations, offering a clearer understanding of how inputs lead to outputs and safeguarding against subtle adversarial attacks [82]. Collectively, these methods make AI systems’ decision-making processes more discernible and reliable, fortifying them against manipulative data inputs.
@@
Table 2: Mechanism/approaches influencing AI safety and the discussion on safety versus transparency, non-discrimination, environmental sustainability and traceability (“-”: negative impact, “+”: positive impact).
@@
Papers Approaches Influencing AI Safety Mitigations towards Potential Attacks Impact Impact on the Other Aspects in AI Regulation
[83] [66]
[67]
@@
Adversarial training Adversarial examples The outcomes of adversarial training yield models with greater semantic significance compared to conventional models. Transparency (+)
[70] [71]
[68]
[69]
@@
Certified robustness Adversarial examples By providing a clear understanding of how the model processes data and makes decisions, thereby allowing for the creation of robustness guarantees and the identification of model vulnerabilities. Transparency (+)
[72] [73]
[74] [75]
[76] [77]
@@
Training data sanitization Data poisoning, backdoor Before initiating machine learning training, ensure the training set is purified by eliminating any tainted samples. Transparency (+)
[78] [79]
[80] [81]
[82]
@@
Robust training Data poisoning, backdoor Adjusting the machine learning training algorithm to conduct resilient training rather than the standard approach. Transparency (+)
[84] [85]
[86] [87]
[88] [89]
[90] [91]
@@
Differential privacy Privacy attacks, training data extraction Enhancing privacy protections for sensitive data used in AI training, testing, and deployment processes. Fairness (+), transparency (+/-), traceability (+)
[92] [93]
[94] [95]
[96] [97]
[98] [99]
@@
Empirical privacy auditing Privacy attacks, training data extraction Empirically assessing an algorithm’s privacy guarantees and establishing lower bounds through privacy attacks. Fairness (+), transparency (+/-), traceability (+)
[100] [101]
[102] [103]
[104]
@@
Machine unlearning Privacy attacks, training data extraction Enhancing user privacy by ensuring that all data that a user wishes to be forgotten from a model is effectively removed. Environmental-sustainability (-), fairness (+)
[105] [106]
[107]
@@
Stricter forward alignment Prompt attacks Developing integrated mechanisms through training for enhanced forward alignment and continuously refining via reinforcement learning informed by human feedback. Fairness (+/-), transparency (+)
@@
[108] Prompt instruction Prompt attacks Prompting the model to handle user input with caution Fairness (+/-)
@@
[105] [108] Stricter backward alignment Prompt attacks Developing built-in safeguards by enhancing backward alignment through training, using tailored benchmark datasets or filters to oversee the input and output of a secure LLM. Transparency (+)
[20] [109]
[41]
@@
Interpretability-based solution Prompt attacks Making the models’ decision-making processes more comprehensible to humans Transparency (+)
@@
Machine unlearning: Machine unlearning represents a novel and distinctive approach aimed at mitigating privacy concerns in machine learning by allowing users to request the removal of their data from trained models, thus potentially enhancing privacy, fairness, and trust in AI systems. This technique comes in two primary forms: exact unlearning, which involves retraining the model from scratch or from a certain checkpoint to ensure complete removal of the data’s influence, and approximate unlearning, where the model’s parameters are adjusted to diminish the influence of the data intended to be forgotten [100][101]. Although promising, the implementation of machine unlearning, whether exact or approximate, introduces complexities related to ensuring the integrity and performance of the model post-unlearning, alongside considerations of computational efficiency and the potential environmental impact of the retraining or updating process [102], [103], [104]. As such, while machine unlearning is a forward step towards user-centric, privacy-preserving AI, it necessitates meticulous design and management to balance its benefits against the operational and ethical challenges it poses.
Differential privacy: By enhancing privacy protections for sensitive data used in AI training, testing, and deployment, Differential Privacy (DP) contributes to the safety of AI systems [84]. DP stands as a robust standard for privacy protection, ensuring that any insights an attacker might gain from an AI system’s output about individual data points are severely limited. The development of this concept was prompted by the discovery of reconstruction attacks, which highlighted the vulnerabilities of aggregate information. Initially, the concept was defined with a privacy parameter, $\epsilon$, which sets a threshold for the likelihood of an attacker discerning whether a specific record is present in the dataset. Over time, this notion has evolved to encompass approximate DP, which introduces an additional parameter, $\delta$, representing the probability of inadvertent information disclosure, alongside $\epsilon$ and Rényi DP [86]. This framework contributes significantly to the safety and security of AI systems by curtailing the risks associated with data exposure.
The primary DP algorithm utilized for training machine learning models is DP-SGD [110], which has seen enhancements in recent times, including DP-FTRL [88] and DP matrix factorization [89]. By definition, DP provides protection against data reconstruction and membership inference attacks. Data reconstruction attacks attempt to reconstruct sensitive information about individuals from the trained model’s outputs, while membership inference attacks aim to determine whether a specific individual’s data was part of the training dataset used to build the model. The definition of DP inherently imposes limits on an adversary’s ability to succeed in mounting membership inference attacks. DP guarantees that the presence or absence of any individual’s data in the training set does not significantly affect the model’s outputs, thus making it difficult for attackers to infer membership with high confidence. Thudi et al. [87] have derived tight bounds on the success of membership inference attacks, further strengthening the efficacy of DP in protecting against such attacks. However, it’s important to note that DP does not provide complete protection against all types of attacks. Specifically, it does not offer guarantees against model extraction attacks, which aim to recover the underlying model itself rather than individual data points. DP is primarily designed to safeguard the privacy of the training data, not the model parameters or architecture. Several papers have highlighted the limitations of using DP to defend against property inference attacks, which seek to extract properties or characteristics of subpopulations within the training dataset [90][91]. While DP can mitigate some aspects of privacy leakage, it may not be sufficient to prevent all forms of property inference attacks.
DP can contribute to fairness by protecting sensitive information and ensuring that individuals from different demographic groups are treated equally in terms of privacy protection [111]. This can help prevent the misuse or discriminatory use of data that could otherwise lead to biased outcomes in decision-making processes. Implementing DP techniques often involves transparent communication about data collection and usage practices [85, 112]. Organizations utilizing DP typically need to be transparent about their methods to protect privacy while still extracting useful insights from data. This transparency can foster trust between organizations and individuals, enhancing overall transparency in data handling practices. However, it can also introduce challenges that may reduce transparency, particularly regarding the complexity of implementation and the trade-offs between privacy and model transparency. DP techniques may enhance traceability by providing mechanisms to track how data is utilized and ensuring that any analysis or inference made from the data preserves individual privacy. By tracking privacy-preserving transformations applied to the data, it becomes possible to trace back the origin of any released information to its source data, thus enabling better accountability and auditing of data usage. While DP can have positive impacts in these areas, it’s essential to note that its effectiveness relies on proper implementation and adherence to best practices. Additionally, there may be trade-offs between privacy and utility in certain applications, where more aggressive privacy protections could limit the usefulness of data for certain analyses. Therefore, careful consideration of the specific context and goals is necessary when applying DP techniques.
Empirical privacy auditing: One of the primary obstacles in implementing DP in practical applications is determining the appropriate privacy parameters to strike a balance between privacy levels and utility, typically assessed through the accuracy of machine learning models [113]. Evaluating privacy-preserving algorithms often relies on worst-case analysis, which may not accurately reflect real-world scenarios, leading to utility loss if parameters are selected solely based on theoretical considerations. As a result, practical applications frequently opt for higher privacy parameters, exemplified by the use of $\epsilon$ 19.61 in the 2020 U.S. Census release, which complicates the precise assessment of the attained level of privacy in real-world scenarios [54].
Lately, an area of new and compelling research direction is the development of privacy auditing, as initiated by Jagielski et al. [92]. This work aims to concretely assess the current levels of privacy security offered by an algorithm. It involves ascertaining the minimum levels of guaranteed privacy by conducting potential threat simulations. One method to carry out such evaluations is through the use of membership inference attacks [93][94]; however, adopting poisoning attacks proves to be considerably more skilled at unearthing the depth of the privacy erosion [92][95], thus providing more noted terms of the power of the particular security offers. Recent progress in privacy auditing has led to improved precision for the Gaussian mechanism [96] and the development of strict statistical techniques that lower the need for large sample sizes by utilizing multiple canaries [97]. Furthermore, two efficient privacy auditing approaches for training a single model have been introduced: one by Steinke et al. [98], which employs various random data canaries without the additional expense of group privacy, and another by Andrew et al.[99], which uses random client canaries along with cosine similarity test statistics to evaluate privacy in federated learning at the user level.

### v2

@@ 2.2.4 Environmental sustainability
In AI regulation, the concept of environmental sustainability emphasizes the creation, implementation, and utilization of AI systems in ways that mitigate adverse environmental effects. This approach recognizes the ecological impact of AI technology, focusing on issues such as energy use, efficient resource management, and promoting ecological sustainability (?). This includes safeguarding the environment and enhancing its quality, which also pertains to human health and safety (?).
Crucial aspects of this environmental focus in AI regulation include promoting energy efficiency, particularly in AI systems with high computational demands like large data centers, which are significant electricity consumers (?). Additionally, it encompasses the pursuit of sustainable development, efforts to decrease the carbon footprint of AI technologies, and the advancement of ‘Green AI’ (?). This latter term refers to designing AI models that are not just high-performing but also optimized for minimal consumption of computational resources and energy. Moreover, it involves adopting eco-friendly practices, such as minimizing electronic waste and encouraging the recycling of AI hardware components.
@@ 2.2.5 Traceability
The principle of traceability within the framework of AI regulation encompasses the systematic capability to monitor, record, and chronicle the decision-making processes employed by AI systems during their entire operational lifespan (?). This principle is crucial for ensuring that AI operations are transparent and accountable. For instance, an AI system should possess robust logging features that provide a traceability level suitable for its intended use, as per the guidelines outlined in the AI Act. Such traceability measures are vital for verifying the AI system’s performance and compliance throughout its lifecycle (?).
Additionally, traceability involves closely tracking the various iterations of AI models, encompassing all modifications and updates that occur over time. This tracking is not just a record-keeping exercise but a critical tool for understanding how an AI model evolves (?). It enables stakeholders to comprehend the changes made to the model, assess their impacts, and, if necessary, revert to prior versions when newer updates do not perform as expected or introduce unforeseen issues.
@@ 2.3 Human-centered AI Act
We categorize stakeholders into four main groups: enterprises, researchers and developers, regulators and policymakers, and users and customers, as shown in Figure 3. According to recent research (?, ?, ?, ?), these stakeholders have distinct interests and responsibilities regarding ethical AI development and deployment. Enterprises are responding to both internal and external pressures to ensure their AI technologies are ethically deployed and do not cause harm. For instance, Microsoft conducted a Human Rights Impact Assessment (HRIA) for AI, and Google has established AI principles that reference human rights (?, ?). These efforts aim to mitigate adverse impacts, ensure corporate social responsibility, and maintain public trust. Regulators and policymakers are responsible for aligning AI with human rights principles throughout each stage of the AI lifecycle. Researchers and developers, who are deeply involved in each lifecycle stage, ensure technical performance and efficiency while promoting responsible AI behavior. They also contribute to the discourse on AI and human rights by examining the social impacts of AI, developing ethical frameworks, and advocating for policies that prioritize human dignity (?). Lastly, end-users and customers focus on the risks and harms associated with AI. They advocate for AI systems that are free from biases, transparent in their operations and decision-making processes, particularly concerning conflicts of interest, and safe and traceable, including maintaining data integrity and security (?).
@@
Ethical concerns in the data preparation stage include ensuring data safety, obtaining informed consent, and addressing biases in data collection (?). It is crucial to collect representative data that accurately reflects the diversity of the population to avoid discrimination and ensure fairness. This process must also ensure the data owners’ safety and make the data collection process transparent and traceable for data providers. During model training, it is essential to build a robust AI against various threats to ensure its safety, while the ethical use of computational resources should be considered to minimize environmental impacts. Transparency and traceability in the deployment process of AI models, including integrating other components within the AI system, are crucial to ensuring user safety. During the inference stage, it is important to evaluate the model’s fairness, environmental sustainability, and safety in real-world scenarios (?), and accordingly can opt to implement bias mitigation, cost-effective, and threats mitigation techniques. The output of the AI system is normally shown as various interactions with customers, AI-enabled decisions and actions, classification labels, prediction values, and/or other content like text, images, etc. All those output data must be transparent and traceable.
@@ 3.1.1 Overview
There’s a concerted effort within the AI security community to compile and comprehend the TTPs that adversaries might deploy against AI systems. A notable contribution in this domain comes from MITRE (?), renowned for their MITRE ATT&CK TTP framework. They have extended their expertise to the AI field by publishing a comprehensive set of TTPs tailored for AI systems (?). This work by MITRE is instrumental in aiding organizations to better understand and prepare for potential threats against their AI infrastructures, reflecting the growing importance of specialized TTPs in the evolving landscape of AI security.
@@
Table 1: Potential Attacks Across Different Stages of the AI Pipeline Ensuring Safety
@@
Data Poisoning Deliberate manipulation of training data to corrupt model learning, causing incorrect predictions. Common methods include injecting malicious data or altering existing data points. Data Acquisition & Preparation (?)(?)
Training Data Extraction Unauthorized extraction or reconstruction of sensitive information from the training dataset used in model development. Model Deployment & Inference (?)
Training-time Attacks in GenAI Attacks targeting pre-training and fine-tuning stages of GenAI models by manipulating portions of training data to cause specific model failures. All stages (?)
Backdoor Attack Insertion of hidden triggers into AI models that cause incorrect behavior only when specific inputs are present, while maintaining normal performance otherwise. All stages (?)(?)
Adversarial Examples Attack Specially crafted inputs that cause AI models to produce incorrect outputs despite appearing normal to humans. Includes white-box, black-box, and transferability attacks. Model Deployment & Inference (?)(?)(?)
Exfiltration Theft of sensitive data, proprietary algorithms, or intellectual property from AI models through unauthorized access or exploitation. Model Deployment & Inference (?)(?)
Inference-time attacks in GenAI Exploitation of vulnerabilities in deployed LLMs and RAG applications during the model’s operational phase. Model Deployment & Inference (?)
Prompt Attacks Manipulation of LLM inputs through carefully crafted prompts to exploit model vulnerabilities and induce unintended behaviors. All stages (?)
Privacy Attacks Unauthorized extraction of sensitive information through techniques like membership inference, model extraction, and property inference attacks. All stages (?)
Ethical Hacking Authorized security testing of AI systems to identify and address vulnerabilities before malicious exploitation. All stages (?)
@@ 3.1.2 Techniques related to safety and their relations to other aspects of the AI Act
Adversarial training: Adversarial training, introduced by Goodfellow et al. (?) and further developed by Madry et al. (?), enhances model safety by incorporating adversarial examples during the training process. This method improves resilience against attacks and increases the model’s semantic content (?), thereby boosting interpretability and predictability. These features are crucial for safety, allowing better anticipation and mitigation of potential failures. However, adversarial training often comes with trade-offs, including reduced accuracy on clean data and increased computational costs.
@@
Formal verification employs methods from formal methods, such as Satisfiability Modulo Theories (SMT) solvers and abstract interpretation, to certify the robustness of neural networks against adversarial inputs (?) (?). This mathematically rigorous approach enhances transparency by providing a solid foundation for understanding and documenting model behavior under adversarial conditions. While formal verification currently faces limitations in scalability and computational efficiency, it offers significant potential for comprehensive robustness certification. This rigorous verification is essential for developing transparent AI systems, as it allows for a deeper understanding of how and why a model is robust against certain types of adversarial inputs.
@@
Certified adversarial robustness: Certified adversarial robustness is crucial for AI system safety and transparency, ensuring predictable and reliable behavior under adversarial inputs. Randomized smoothing enhances classifier robustness by integrating Gaussian noise into input data (?)(?), generating predictions most likely correct under noise perturbations. This approach yields provably robust classifiers, particularly against $\ell_{2}$-norm attacks (?), enhancing system predictability and transparency. The method offers certified robustness for a subset of testing samples, providing quantifiable metrics to track model performance under adversarial conditions. Recent advancements have extended this concept to more complex perturbations, integrating denoising diffusion probabilistic models with high-accuracy classifiers, broadening the scope of transparency across various adversarial scenarios (?).
@@
Training data sanitization: Training data sanitization enhances AI system safety by mitigating poisoned samples’ impact. These methods identify and remove malicious samples before training. Techniques like the Region of Non-Interest (RONI) method enhance model reliability by excluding poisoned samples (?). Label cleaning methods target label flipping attacks (?), while outlier detection (?) and clustering methods (?)(?) remove anomalies. Ensemble variance computation helps identify potential poisoning attempts (?). Cybersecurity mechanisms for provenance and integrity attestation add further protection. Overall, data sanitization improves AI transparency by preserving training data authenticity, enhancing system safety and decision-making reliability. Robust training: Robust training enhances the transparency and safety of AI systems by integrating techniques like ensemble model voting, robust optimization, and randomized smoothing (?)(?). Ensemble models allow for comparative analysis of individual model decisions, improving insight into the decision-making process. Robust optimization, through methods like trimmed loss functions, minimizes the impact of extreme data points, making the model’s reasoning more transparency and less prone to data poisoning (?)(?). Randomized smoothing adds stability to model predictions against minor input perturbations, offering a clearer understanding of how inputs lead to outputs and safeguarding against subtle adversarial attacks (?). Collectively, these methods make AI systems’ decision-making processes more discernible and reliable, fortifying them against manipulative data inputs.
@@
Table 2: Mechanism/approaches influencing AI safety and the discussion on safety versus other four aspects(“-”: negative impact, “+”: positive impact).
@@
Papers Approaches Influencing AI Safety Mitigations towards Potential Attacks Impact Other Aspects’ Impacts in AI Regulation
@@
(?, ?, ?) Adversarial training Adversarial examples The outcomes of adversarial training yield models with greater semantic significance compared to conventional models. Transparency (+)
@@
(?, ?, ?, ?) Certified robustness Adversarial examples By providing a clear understanding of how the model processes data and makes decisions, thereby allowing for the creation of robustness guarantees and the identification of model vulnerabilities. Transparency (+)
@@
(?, ?, ?, ?) Training data sanitization Data poisoning, backdoor Before initiating machine learning training, ensure the training set is purified by eliminating any tainted samples. Transparency (+)
@@
(?, ?, ?, ?) Robust training Data poisoning, backdoor Adjusting the machine learning training algorithm to conduct resilient training rather than the standard approach. Transparency (+)
@@
(?, ?, ?) Differential privacy Privacy attacks, training data extraction Enhancing privacy protections for sensitive data used in AI training, testing, and deployment processes. Fairness (+), transparency (+/-), traceability (+)
@@
(?, ?, ?) Empirical privacy auditing Privacy attacks, training data extraction Empirically assessing an algorithm’s privacy guarantees and establishing lower bounds through privacy attacks. Fairness (+), transparency (+/-), traceability (+)
@@
(?, ?) Machine unlearning Privacy attacks, training data extraction Enhancing user privacy by ensuring that all data that a user wishes to be forgotten from a model is effectively removed. Environmental-sustainability (-), fairness (+)
@@
(?, ?) Stricter forward alignment Prompt attacks Developing integrated mechanisms through training for enhanced forward alignment and continuously refining via reinforcement learning informed by human feedback. Fairness (+/-), transparency (+)
@@
(?) Prompt instruction Prompt attacks Prompting the model to handle user input with caution Fairness (+/-)
@@
(?, ?) Stricter backward alignment Prompt attacks Developing built-in safeguards by enhancing backward alignment through training, using tailored benchmark datasets or filters to oversee the input and output of a secure LLM. Transparency (+)
@@
(?, ?, ?) Interpretability-based solution Prompt attacks Making the models’ decision-making processes more comprehensible to humans Transparency (+)
@@
Machine unlearning: Machine unlearning represents a novel and distinctive approach aimed at mitigating privacy concerns in machine learning by allowing users to request the removal of their data from trained models, thus potentially enhancing privacy, fairness, and trust in AI systems. This technique comes in two primary forms: exact unlearning, which involves retraining the model from scratch or from a certain checkpoint to ensure complete removal of the data’s influence, and approximate unlearning, where the model’s parameters are adjusted to diminish the influence of the data intended to be forgotten (?)(?). Although promising, the implementation of machine unlearning, whether exact or approximate, introduces complexities related to ensuring the integrity and performance of the model post-unlearning, alongside considerations of computational efficiency and the potential environmental impact of the retraining or updating process (?), (?), (?). As such, while machine unlearning is a forward step towards user-centric, privacy-preserving AI, it necessitates meticulous design and management to balance its benefits against the operational and ethical challenges it poses.
Differential privacy: Differential Privacy (DP) enhances AI system safety by protecting sensitive training data (?). It limits an attacker’s ability to gain insights about individual data points from AI outputs. DP has evolved to include approximate DP and Rényi DP (?). The primary algorithm for training machine learning models is DP-SGD (?), with recent enhancements like DP-FTRL (?) and DP matrix factorization (?). DP protects against data reconstruction and membership inference attacks, with tight bounds derived by Thudi et al. (?). However, it doesn’t fully protect against model extraction or property inference attacks (?)(?).
DP contributes to fairness by ensuring equal privacy protection across demographic groups (?). It enhances transparency in data handling practices (?, ?) and improves traceability of data usage. However, trade-offs between privacy and utility must be carefully considered in specific applications.
Empirical privacy auditing: Implementing Differential Privacy (DP) faces challenges in balancing privacy and utility (?). Worst-case analysis often leads to higher privacy parameters in practice, as seen in the 2020 U.S. Census (?). Privacy auditing, initiated by Jagielski et al. (?), aims to assess actual privacy levels through threat simulations. While membership inference attacks (?)(?) are used, poisoning attacks are more effective at revealing privacy erosion (?)(?).
Recent advancements include improved precision for the Gaussian mechanism (?) and techniques reducing large sample size requirements (?). Efficient privacy auditing approaches have been introduced by Steinke et al. (?) using random data canaries, and Andrew et al. (?) employing random client canaries for federated learning privacy evaluation.